a biologist estimates that 50 of deer in the region carry a certain type of tick. For a sample of 300 deer selected at random, what is the chance that 155 or fewer deer have this tick?


1
Expert's answer
2021-05-11T14:06:54-0400

Let X denote number of deers in the region carry a certain type of tick


Here X follows binomial distribution with parameters "n=300 \text{ and } p=0.16"

From the normal approximation to the binomial variable-


"z=\dfrac{x-np}{\sqrt{np(1-p)}}" is the standard normal variable.


Probability that 155 or fewer deer have the tick-


"P(X\le 155)=P(X\le 155+0.5)=P(X<155.5)"


 "=P(z\le \dfrac{155.5-(300)(0.16)}{\sqrt{300(0.16(1-0.16)}})\\[9pt]=P(z\le \dfrac{107.5}{6.349})\\[9pt]=P(z\le 16.93)"


   "=P(Z\le 16.93)=0.7782"